OpenStudy (anonymous):

square root of 81a^10b^6

OpenStudy (anonymous):

Recall that the square root of a variable is equal to halving the exponent. So, the square root of a^10 is a^(10/2) = a^5 The square root of b^6 is b^(6/2)=b^3 The square root of 81 is 9. So our answer is \[9a^5 b^3\]